Certain number of men can finish a work in 10 days. If however there were 10 men less it then it will take 10 days more for work to be finished. How many men were there originally?
1
Expert's answer
2011-04-13T11:34:03-0400
Let's denote amount of work as X, the number of men as Y, thus 1 man can do in one day& X/(10Y) of the total work. If there were 10 men less (Y-10) it will take 20 days to do all the work:
(Y-10)*20*& X/(10Y) =& X;
thus
(Y-10)/Y = 1/2
Y = 20.
Answer: there were 20 men originally.